OpenStudy (anonymous):

Find the value that makes the ratio equal 98;46,49: ?

OpenStudy (anonymous):

23

OpenStudy (anonymous):

i think its 1

OpenStudy (anonymous):

sorry, here's how >.< 98/46 = 49/n 98n = 2254 so n = 23

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

welcome [: